# Bridging Gulfs in UI Generation through Semantic Guidance

### v1

@@ Abstract.
While generative AI enables high-fidelity UI generation from text prompts, users struggle to articulate design intent and evaluate or refine results—creating gulfs of execution and evaluation. To understand the information needed for UI generation, we conducted a thematic analysis of UI prompting guidelines, identifying key design semantics and discovering that they are hierarchical and interdependent. Leveraging these findings, we developed a system that enables users to specify semantics, visualize relationships, and extract how semantics are reflected in generated UIs. By making semantics serve as an intermediate representation between human intent and AI output, our system bridges both gulfs by making requirements explicit and outcomes interpretable. A comparative user study suggests that our approach enhances users’ perceived control over intent expression, outcome interpretation, and facilitates more predictable, iterative refinement. Our work demonstrates how explicit semantic representation enables systematic and explainable exploration of design possibilities in AI-driven UI design.
@@ 1. Introduction
The rapid advancement of generative AI has introduced new possibilities for the landscape of user interface (UI) design (Nielsen, 2024). Today, users can generate high-fidelity user interfaces through natural language prompts, democratizing access to sophisticated design capabilities (Chen et al., 2025). However, this apparent simplicity masks significant challenges in bidirectional communication between users and AI systems: users struggle to effectively express their design intentions, interpret generated outputs, and predict the effects of modifications.
@@
These challenges reflect deeper issues identified in interaction design theory as the gulf of execution (the distance between user intentions and available actions) and the gulf of evaluation (the effort required to interpret system state and determine if goals were achieved) (Hutchins et al., 1986). In generative UI systems, both gulfs are particularly wide: users cannot easily express their design intentions through prompts (execution), nor can they readily understand how their prompts influenced the generated output (evaluation), and they amplify each other through iteration.
To address these challenges, we conducted a systematic investigation into how design intent can be effectively communicated to and interpreted by generative AI systems. Our approach began with a thematic analysis of prompting guidelines from leading UI generation services, through which we identified and categorized the semantic elements essential for UI generation. This analysis revealed that UI generation semantics operate across multiple levels of abstraction, from high-level product vision to specific component properties. Building on this framework, we developed a semantic-based UI generation system that bridges the gulfs between user intent and AI output. Rather than treating prompts as direct instructions, our approach introduces an intermediate semantic layer that serves as a bridge between natural language and generated designs. This semantics enables bidirectional interaction—both parsing user intentions into structured representations and extracting implemented semantics from generated outputs. We expose the intermediate semantic layer that connects what users want to what the system generates, allowing users to see and control how their design intentions are composed and translated into actual UI elements. Furthermore, the system reveals semantic relationships, helping users identify conflicts and maintain coherence. This transparency enables what was previously a black-box operation into an inspectable and controllable pipeline.
@@ 2.2. Design Intent Specification in Generative AI
As UI generation shifts to natural-language prompts, practitioners face a significant articulation gap in conveying their intent. Chen et al. (Chen et al., 2025) observed that UX practitioners iteratively trial-and-error with prompts, spending considerable time reformulating descriptions, and still frequently end up with UIs that miss their envisioned goals.
@@
Within UI work specifically, researchers also target the evaluation and alignment of intent with generated artifacts. Duan et al. (Duan et al., 2024) provide actionable automatic feedback on mockups using LLM. At the workflow level, studies find UX designers currently see generative AI as assistive rather than substitutive and that those tools can both reduce and introduce collaboration conflicts depending on the quality and transparency of outputs (Li et al., 2024). Bridging requirements and prototypes, Kolthoff et al. (Kolthoff et al., 2024) semi-automatically link user stories to GUI elements, detecting unimplemented stories and recommending components. Across these strands, the through-line is clear: effective intent specification benefits from structured, inspectable intermediates rather than raw text alone. However, existing approaches typically address either input structuring or output evaluation, and most operate on general-purpose dimensions rather than UI-specific semantics. Our approach contributes a domain-specific semantic framework derived from multifaceted UI prompting practices, combined with bidirectional semantic guidance that both parse user intent into a structured form and extract how semantics are realized in generated outputs—connecting specification and evaluation within a unified intermediate layer.
@@ 2.3. UI Semantic Representations
More recent data-driven work has developed representations that encode UIs in terms of their constituent elements and relationships, enabling computational analysis and generation. The Rico dataset (Deka et al., 2017; Liu et al., 2018) made this concrete at scale with screens with view hierarchies, text, and interaction traces, seeding work that treats a UI as a composition of elements and constraints rather than a bitmap. Building atop such corpora, Screen2Vec (Li et al., 2021) learns embeddings from screens and components from hierarchy context and usage. Complementing embedding approaches, Graph4GUI (Jiang et al., 2024) represents each UI as a graph that captures visuo-spatial and semantic relations between elements, improving layout understanding with alignment and containment constraints. Parallel efforts align vision and language semantics to make quality and intent measurable: UIClip (Wu et al., 2024) scores a UI screenshot against a natural-language description, providing a data-driven signal of design quality/fit that can steer generation and exemplar search. Recent UI-tuned VLMs, such as ILuvUI (Jiang et al., 2025), instruction-tuned models to suggest descriptions, QA, and multi-step planning, broadening the kinds of semantic questions one can ask of an interface.
Beyond single screens, newer work elevates task-level semantics so that intent and behavior remain traceable across flows. Park et al. (Park et al., 2023) formulates app-level retrieval and assesses design consistency. S&UI (Park et al., 2025) extracts multiple level UI semantics for UI inspiration search, indicating that explicit semantics help designers find and reason about options at the level they think. Recent work has also explored intermediate representations for generative UIs, including task-driven data models (Cao et al., 2025) and malleable overview-detail interfaces (Min et al., 2025; Min and Xia, 2025). While these advances address task structure and pattern-level customization, they do not explicitly model the hierarchical relationships between design intents across abstraction levels. Our approach addresses this gap from a complementary angle: by analyzing prompting guidelines from existing UI generation services, we derive a semantic framework that structures hierarchical, interdependent information for generation, providing an inspectable intermediate that links designer intent to concrete UI decisions.
@@ 3. Gulfs in Current UI Generation
While generative AI has significantly advanced capabilities, transforming UI creation from direct manipulation to natural language interaction, this shift has introduced fundamental challenges that mirror and extend Norman’s gulfs of execution and evaluation (Hutchins et al., 1986; Subramonyam et al., 2024). Through the analysis of a recent comprehensive study on UI generation usage (Chen et al., 2025), we identify how these interaction challenges manifest themselves in text-to-UI systems and, critically, how they amplify each other through iterative refinement.
@@ 3.1. Gulf of Execution:
The gulf of execution in UI generation manifests as designers’ inability to determine what to express and how to articulate it in prompts. Unlike traditional design tools (e.g., Figma (Figma, 2025b)), where actions directly manipulate visual elements, generative systems require translating multidimensional design intentions—visual aesthetics, functional requirements, and user contexts—into linear text sequences. Users face two fundamental uncertainties: identifying which design dimensions require explicit specification, and finding a suitable way to prompt concepts that often lack precise verbal equivalents. For instance, when designing an e-commerce checkout page, the user must decide which elements to specify, how visually appealing they are, and how to induce purchase. Even when users know what they want, conceptual descriptors such as “modern,” “professional,” “trustworthy” carry vastly different interpretations: should “trustworthy” translate to blue colors (may be associated with trust), serif typography (traditional authority), or increased whitespace (clarity)?
@@ 3.2. Gulf of Evaluation:
The gulf of evaluation emerges when designers cannot understand how their inputs are translated into specific design decisions or assess whether the generated output appropriately reflects their intentions. When AI systems produce complete interfaces from brief prompts, they make numerous implicit decisions about layout, colors, and interaction patterns, none traceable to specific prompt elements. For example, when a “professional dashboard” prompt could yield a dark theme, users cannot determine whether this reflects the AI’s interpretation of “professional,” a common dashboard pattern, or an arbitrary aesthetic choice. This opacity transforms evaluation from a deliberate assessment process into detective work, where users have to reverse-engineer the AI’s interpretation. The challenge intensifies when outputs appear visually polished but may not align with unstated functional requirements or design principles, making it difficult to distinguish between acceptable alternatives and fundamental misunderstandings of intent.
@@ 4. Understanding UI Generation Semantics
Figure 2. From prompting guidelines to a semantic framework. We conducted a thematic analysis of prompting guidelines from major UI generation tools, surfacing recurring patterns of what information users specify. This yields a four-level hierarchical representation — Refer to caption Product, Refer to caption Design System, Refer to caption Feature, Refer to caption Component — that organizes interdependent design semantics from high- to low-levels. Semantic elements are related both vertically (between levels) and horizontally (within levels), meaning that changes to one element can cascade across others.
@@ 4.1. Analysis Process
We collected prompting guidelines with example prompts, which contain best practice recommendations, from six leading prompt-based UI generation services (Vercel v0 (Vercel, 2025), Google Stitch (Labs, 2025), Figma Make (Figma, 2025a), Uizard (Technologies, 2024), Lovable (Lovable, 2023), and Relume (Relume, 2025)), resulting in 907 guide fragments consisting of 601 guideline and 306 prompt examples. Three authors collaboratively analyzed these materials through iterative coding, identifying recurring patterns and categorizing the types of information required for different UI generation tasks. We conducted exploratory analysis, allowing patterns to emerge from the data while maintaining a focus on what information users need to provide and how it relates to generation outcomes.
@@ 4.2. Hierarchical Semantic Framework
Our analysis revealed that UI generation encompasses multiple task types and scopes. We identified three primary action types: creation (generating new UI from scratch), modification (altering existing designs), and analysis (extracting or evaluating design properties). These actions operate at different scope levels: entire applications, partial sections or screens, and individual elements. Each combination of action type and scope requires different semantic information, and effective UI generation must coordinate this information across multiple abstraction levels, where changes at one level can influence other levels. Across these varying tasks and scopes, the semantic information organizes into four hierarchical levels: [Uncaptioned image] Product[Uncaptioned image] Design System[Uncaptioned image] Feature, and  [Uncaptioned image] Component. While this structure resembles Garrett’s layer of user experience (Garrett, 2010) and the visual-building logic of Atomic Design (Frost, 2016), our framework spans both, from high-level experience to specific elements. In doing so, it captures the broad semantics expressed in generative workflows and the relationships that shape how those semantics influence generated outcomes.
@@
To illustrate the hierarchical relationship of these semantics, consider a music streaming product such as Spotify. At the [Uncaptioned image] Product level, defining the Goal as “providing a personalized and immersive audio experience” for a Target User who values “effortless connection to music” sets the foundational context. This intent relates to the [Uncaptioned image] Design System: to support immersion, the Color is defined as “Dark Mode” to make album artwork serve as the primary visual focus, while the Visual Mood is set to “energetic and bold” to guide stylistic choices. Descending to the [Uncaptioned image] Feature level, exemplified by the home screen of the music app, the Function of “personalized recommendation hub” drives the Information Architecture to organize content into “horizontal scrollable feeds” rather than vertical lists, prioritizing visual scanning over textual density. Finally, at the [Uncaptioned image] Component level, these constraints materialize in specific Interactivity and Properties: an Album Card is implemented not merely as an image, but with a “hover-triggered Play button” and “rounded corners,” serving the high-level product goal of reducing friction for playback.
@@ 4.3. Key Insights
As illustrated in the walkthrough, a critical characteristic of our framework analysis is the relationships between semantic elements. These relationships operate both vertically (between levels) and horizontally (within levels). Understanding these relationships is essential because changes to one semantic element can have far-reaching effects throughout the design in the generation context. For example, changing the target user from “young professionals” to “older adult users” doesn’t just affect font size; it influences color contrast, interaction patterns, information density, and navigation complexity. Together, these patterns indicate that prompt-based UI generation requires coordinating meaning across multiple abstraction layers rather than simply describing visual appearance—helping explain why current text-based prompting feels difficult and unpredictable. Users must somehow encode this multi-level, interconnected information into linear text, without clear guidance on what to specify explicitly versus what to leave for UI generation.
@@ 5. System
To bridge the gulfs identified in Section 3, we designed and implemented a system that leverages the semantic framework from Section 4 as an intermediate representation between human intent and AI-generated user interface. We developed a web-based application that provides structured mechanisms for specifying, analyzing, and refining UI designs through explicit semantic manipulation.
@@ 2. DG2
    Transparent Analysis tackles the gulf of evaluation by extracting and presenting the semantic decisions that the model enacted in the generated UI. Instead of treating AI generation as a black box, the system analyzes the output to reveal what design choices were made and how they relate to user input. This transparency enables users to understand how their intents were actually generated.
@@ 3. DG3
Figure 4. Example of input semantics specification. Users can define semantics manually or via natural language prompts. Natural language input is Refer to caption parsed into structured semantics across our four-level framework: Refer to caption Product—Description, Refer to caption Design System—Design Style, Color, Refer to caption Feature—Function, and Refer to caption Component—Card Component.

### v2

@@ Abstract.
While generative AI enables high-fidelity UI generation from text prompts, users struggle to articulate design intent and evaluate or refine results—creating gulfs of execution and evaluation. To understand the information needed for UI generation, we conducted a thematic analysis of UI prompting guidelines, identifying key design semantics and discovering that they are hierarchical and interdependent. Leveraging these findings, we developed a system that enables users to specify semantics, visualize relationships, and extract how semantics are reflected in generated UIs. By making semantics serve as an intermediate representation between human intent and AI output, our system bridges both gulfs by making requirements explicit and outcomes interpretable. A comparative user study suggests that our approach enhances users’ perceived control over intent expression and outcome interpretation, and facilitates more predictable iterative refinement. Our work demonstrates how explicit semantic representation enables systematic and explainable exploration of design possibilities in AI-driven UI design.
@@ 1. Introduction
The rapid advancement of generative AI has introduced new possibilities in the landscape of user interface (UI) design (Nielsen, 2023). Today, users can generate high-fidelity user interfaces through natural language prompts, democratizing access to sophisticated design capabilities (Chen et al., 2025). However, this apparent simplicity masks significant challenges in bidirectional communication between users and AI systems: users struggle to effectively express their design intentions, interpret generated outputs, and predict the effects of modifications.
@@
These challenges reflect deeper issues identified in interaction design theory as the gulf of execution (the distance between user intentions and available actions) and the gulf of evaluation (the effort required to interpret system state and determine if goals were achieved) (Hutchins et al., 1985). In generative UI systems, both gulfs are particularly wide: users cannot easily express their design intentions through prompts (execution), nor can they readily understand how their prompts influenced the generated output (evaluation). Moreover, these gulfs amplify each other through iteration.
To address these challenges, we conducted a systematic investigation into how design intent can be effectively communicated to and interpreted by generative AI systems. Our approach began with a thematic analysis of prompting guidelines from leading UI generation services, through which we identified and categorized the semantic elements essential for UI generation. This analysis revealed that UI generation semantics operate across multiple levels of abstraction, from high-level product vision to specific component properties. Building on this framework, we developed a semantic-based UI generation system that bridges the gulfs between user intent and AI output. Rather than treating prompts as direct instructions, our approach introduces an intermediate semantic layer that serves as a bridge between natural language and generated designs. These semantics enable bidirectional interaction—both parsing user intentions into structured representations and extracting implemented semantics from generated outputs. We expose the intermediate semantic layer that connects what users want to what the system generates, allowing users to see and control how their design intentions are composed and translated into actual UI elements. Furthermore, the system reveals semantic relationships, helping users identify conflicts and maintain coherence. This transparency enables what was previously a black-box operation to become an inspectable and controllable pipeline.
@@ 2.2. Design Intent Specification in Generative AI
As UI generation shifts to natural-language prompts, practitioners face a significant articulation gap in conveying their intent. Chen et al. (Chen et al., 2025) observed that UX practitioners iteratively engage in trial-and-error with prompts, spending considerable time reformulating descriptions, and still frequently end up with UIs that miss their envisioned goals.
@@
Within UI work specifically, researchers also target the evaluation and alignment of intent with generated artifacts. Duan et al. (Duan et al., 2024) provide actionable automatic feedback on mockups using LLMs. At the workflow level, studies find UX designers currently see generative AI as assistive rather than substitutive and that those tools can both reduce and introduce collaboration conflicts depending on the quality and transparency of outputs (Li et al., 2024). Bridging requirements and prototypes, Kolthoff et al. (Kolthoff et al., 2024) semi-automatically link user stories to GUI elements, detecting unimplemented stories and recommending components. Across these strands, the through-line is clear: effective intent specification benefits from structured, inspectable intermediates rather than raw text alone. However, existing approaches typically address either input structuring or output evaluation, and most operate on general-purpose dimensions rather than UI-specific semantics. Our approach contributes a domain-specific semantic framework derived from multifaceted UI prompting practices, combined with bidirectional semantic guidance that both parses user intent into a structured form and extracts how semantics are realized in generated outputs—connecting specification and evaluation within a unified intermediate layer.
@@ 2.3. UI Semantic Representations
More recent data-driven work has developed representations that encode UIs in terms of their constituent elements and relationships, enabling computational analysis and generation. The Rico dataset (Deka et al., 2017; Liu et al., 2018) made this concrete at scale with screens containing view hierarchies, text, and interaction traces, seeding work that treats a UI as a composition of elements and constraints rather than a bitmap. Building atop such corpora, Screen2Vec (Li et al., 2021) learns embeddings from screens and components from hierarchy context and usage. Complementing embedding approaches, Graph4GUI (Jiang et al., 2024) represents each UI as a graph that captures visuo-spatial and semantic relations between elements, improving layout understanding with alignment and containment constraints. Parallel efforts align vision and language semantics to make quality and intent measurable: UIClip (Wu et al., 2024) scores a UI screenshot against a natural-language description, providing a data-driven signal of design quality/fit that can steer generation and exemplar search. Recent fine-tuned VLMs, such as ILuvUI (Jiang et al., 2025), have been instruction-tuned to suggest descriptions, QA, and multi-step planning, broadening the kinds of semantic questions one can ask of an interface.
Beyond single screens, newer work elevates task-level semantics so that intent and behavior remain traceable across flows. Park et al. (Park et al., 2023) formulate app-level retrieval and assess design consistency. S&UI (Park et al., 2025) extracts multi-level UI semantics for UI inspiration search, indicating that explicit semantics help designers find and reason about options at the level they think. Recent work has also explored intermediate representations for generative UIs, including task-driven data models (Cao et al., 2025) and malleable overview-detail interfaces (Min et al., 2025; Min and Xia, 2025). While these advances address task structure and pattern-level customization, they do not explicitly model the hierarchical relationships between design intents across abstraction levels. Our approach addresses this gap from a complementary angle: by analyzing prompting guidelines from existing UI generation services, we derive a semantic framework that structures hierarchical, interdependent information for generation, providing an inspectable intermediate that links designer intent to concrete UI decisions.
@@ 3. Gulfs in Current UI Generation
While generative AI has significantly advanced capabilities, transforming UI creation from direct manipulation to natural language interaction, this shift has introduced fundamental challenges that mirror and extend Norman’s gulfs of execution and evaluation (Hutchins et al., 1985; Subramonyam et al., 2024). Through the analysis of a recent comprehensive study on UI generation usage (Chen et al., 2025), we identify how these interaction challenges manifest themselves in text-to-UI systems and, critically, how they amplify each other through iterative refinement.
@@ 3.1. Gulf of Execution:
The gulf of execution in UI generation manifests as designers’ inability to determine what to express and how to articulate it in prompts. Unlike traditional design tools (e.g., Figma (Figma, 2025b)), where actions directly manipulate visual elements, generative systems require translating multidimensional design intentions—visual aesthetics, functional requirements, and user contexts—into linear text sequences. Users face two fundamental uncertainties: identifying which design dimensions require explicit specification, and finding a suitable way to prompt concepts that often lack precise verbal equivalents. For instance, when designing an e-commerce checkout page, the user must decide which elements to specify, how visually appealing they should be, and how to induce purchase. Even when users know what they want, conceptual descriptors such as “modern,” “professional,” and “trustworthy” carry vastly different interpretations: should “trustworthy” translate to blue colors (associated with trust), serif typography (traditional authority), or increased whitespace (clarity)?
@@ 3.2. Gulf of Evaluation:
The gulf of evaluation emerges when designers cannot understand how their inputs are translated into specific design decisions or assess whether the generated output appropriately reflects their intentions. When AI systems produce complete interfaces from brief prompts, they make numerous implicit decisions about layout, colors, and interaction patterns, none traceable to specific prompt elements. For example, if a “professional dashboard” prompt yields a dark theme, users cannot determine whether this reflects the AI’s interpretation of “professional,” a common dashboard pattern, or an arbitrary aesthetic choice. This opacity transforms evaluation from a deliberate assessment process into detective work, where users have to reverse-engineer the AI’s interpretation. The challenge intensifies when outputs appear visually polished but may not align with unstated functional requirements or design principles, making it difficult to distinguish between acceptable alternatives and fundamental misunderstandings of intent.
@@ 4. Understanding UI Generation Semantics
Figure 2. From prompting guidelines to a semantic framework. We conducted a thematic analysis of prompting guidelines from major UI generation tools, surfacing recurring patterns of what information users specify. This yields a four-level hierarchical representation — Refer to caption Product, Refer to caption Design System, Refer to caption Feature, Refer to caption Component — that organizes interdependent design semantics from high to low levels. Semantic elements are related both vertically (between levels) and horizontally (within levels), meaning that changes to one element can cascade across others.
@@ 4.1. Analysis Process
We collected prompting guidelines with example prompts, which contain best practice recommendations, from six leading prompt-based UI generation services (Vercel v0 (Vercel, 2025), Google Stitch (Labs, 2025), Figma Make (Figma, 2025a), Uizard (Technologies, 2024), Lovable (Lovable, 2023), and Relume (Relume, 2025)), resulting in 907 guide fragments consisting of 601 guidelines and 306 prompt examples. Three authors collaboratively analyzed these materials through iterative coding, identifying recurring patterns and categorizing the types of information required for different UI generation tasks. We conducted exploratory analysis, allowing patterns to emerge from the data while maintaining a focus on what information users need to provide and how it relates to generation outcomes.
@@ 4.2. Hierarchical Semantic Framework
Our analysis revealed that UI generation encompasses multiple task types and scopes. We identified three primary action types: creation (generating new UI from scratch), modification (altering existing designs), and analysis (extracting or evaluating design properties). These actions operate at different scope levels: entire applications, partial sections or screens, and individual elements. Each combination of action type and scope requires different semantic information, and effective UI generation must coordinate this information across multiple abstraction levels, where changes at one level can influence other levels. Across these varying tasks and scopes, the semantic information organizes into four hierarchical levels: [Uncaptioned image] Product[Uncaptioned image] Design System[Uncaptioned image] Feature, and  [Uncaptioned image] Component. While this structure resembles Garrett’s layers of user experience (Garrett, 2010) and the visual-building logic of Atomic Design (Frost, 2016), our framework encompasses both perspectives, spanning from high-level experience to specific elements. In doing so, it captures the broad semantics expressed in generative workflows and the relationships that shape how those semantics influence generated outcomes.
@@
To illustrate the hierarchical relationships of these semantics, consider a music streaming product such as Spotify. At the [Uncaptioned image] Product level, defining the Goal as “providing a personalized and immersive audio experience” for a Target User who values “effortless connection to music” sets the foundational context. This intent relates to the [Uncaptioned image] Design System: to support immersion, the Color is defined as “Dark Mode” to make album artwork serve as the primary visual focus, while the Visual Mood is set to “energetic and bold” to guide stylistic choices. Descending to the [Uncaptioned image] Feature level, exemplified by the home screen of the music app, the Function of “personalized recommendation hub” drives the Information Architecture to organize content into “horizontal scrollable feeds” rather than vertical lists, prioritizing visual scanning over textual density. Finally, at the [Uncaptioned image] Component level, these constraints materialize in specific Interactivity and Properties: an Album Card is implemented not merely as an image, but with a “hover-triggered Play button” and “rounded corners,” serving the high-level product goal of reducing friction for playback.
@@ 4.3. Key Insights
As illustrated in the walkthrough, critical characteristics of our framework analysis are the relationships between semantic elements. These relationships operate both vertically (between levels) and horizontally (within levels). Understanding these relationships is essential because changes to one semantic element can have far-reaching effects throughout the design in the generation context. For example, changing the target user from “young professionals” to “older adult users” doesn’t just affect font size; it influences color contrast, interaction patterns, information density, and navigation complexity. Together, these patterns indicate that prompt-based UI generation requires coordinating meaning across multiple abstraction layers rather than simply describing visual appearance—helping explain why current text-based prompting feels difficult and unpredictable. Users must somehow encode this multi-level, interconnected information into linear text, without clear guidance on what to specify explicitly versus what to leave for UI generation.
@@ 5. System
To bridge the gulfs identified in Section 3, we designed and implemented a system that leverages the semantic framework from Section 4 as an intermediate representation between human intent and AI-generated user interfaces. We developed a web-based application that provides structured mechanisms for specifying, analyzing, and refining UI designs through explicit semantic manipulation.
@@ 2. DG2
    Transparent Analysis tackles the gulf of evaluation by extracting and presenting the semantic decisions that the model enacted in the generated UI. Instead of treating AI generation as a black box, the system analyzes the output to reveal what design choices were made and how they relate to user input. This transparency enables users to understand how their intents were actually implemented.
@@ 3. DG3
Figure 4. Example of input semantics specification. Users can define semantics manually or via natural language prompts. Natural language input is Refer to caption parsed into structured semantics across our four-level framework: Refer to caption Product—Description, Refer to caption Design System—Design Style and Color, Refer to caption Feature—Function, and Refer to caption Component—Card Component.